OpenStudy (anonymous):

what is the difference of 7 6/8 - 3 4/8 =

OpenStudy (anonymous):

4 and 1/4

OpenStudy (anonymous):

Turning, 6/8 - 4/8 = 2/8 or 1/4

OpenStudy (anonymous):

Do decimals, 7.75 - 3.5 = 4.25 or 4 and 1/4

OpenStudy (turingtest):

you could do it like 7+6/8-3-4/8=4+2/8=4 1/4 hmm... what did I do earlier?

OpenStudy (turingtest):

oh I just can't subtract is all LOL

OpenStudy (anonymous):

Not sure, but as long as we have the correct answer it doesn't matter. I didn't look at your equatio too long. lol

OpenStudy (turingtest):

62/8-28/8=36/8 is what I did wrong

OpenStudy (anonymous):

LOL at least we caught it. Thanks for being helpful to everyone tho!